Can someone help me solve this problem?
I need to complete the ordered pairs so that each is a solution for the given equation: 3x + 4y = 12
pair 1 is (0,)
Pair 2 is (, 3/4)
pair 3 is (,0)
pair 4 is (8/3,)
Thanks,
Sher

Found 2 solutions by checkley71, tiffanyc:
Answer by checkley71(8403)

(0,3) just substitute the given value x=0 & solve for y
(3,3/4) just substitute the given value y=3/4 & solve for x
(4,0)
(8/3,1)

Answer by tiffanyc(9)

The coordinates you are given are the x and y values for the linear equation. Simply substitue the coordinate you are given into the equation and isolate the other variable:
Pair 1 (0,) so x=0 and we try to find y
3(0)+4y=12.......4y=12...........y=3
so the first pair is (0,3)
Pair 2 (,3/4) so y=3/4 and we try to find x
3x+ 4(3/4)=12 ............3x+3=12.........3x=9.....x=3

And so on and so on